# A Survey on Deep Learning Hardware Accelerators for Heterogeneous HPC Platforms

### v2

@@ Abstract.
Recent trends in deep learning (DL) imposed hardware accelerators as the most viable solution for several classes of high-performance computing (HPC) applications such as image classification, computer vision, and speech recognition. This survey summarizes and classifies the most recent advances in designing DL accelerators suitable to reach the performance requirements of HPC applications. In particular, it highlights the most advanced approaches to support deep learning accelerations including not only GPU and TPU-based accelerators but also design-specific hardware accelerators such as FPGA-based and ASIC-based accelerators, Neural Processing Units, open hardware RISC-V-based accelerators and co-processors. The survey also describes accelerators based on emerging memory technologies and computing paradigms, such as 3D-stacked Processor-In-Memory, non-volatile memories (mainly, Resistive RAM and Phase Change Memories) to implement in-memory computing, Neuromorphic Processing Units, and accelerators based on Multi-Chip Modules. Among emerging technologies, we also include some insights into quantum-based accelerators and photonics. To conclude, the survey classifies the most influential architectures and technologies proposed in the last years, with the purpose of offering the reader a comprehensive perspective in the rapidly evolving field of deep learning.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXjournal: CSURccs: Computer systems organization Architecturesccs: Hardware Reconfigurable logic and FPGAsccs: Hardware Emerging technologiesccs: Hardware Very large scale integration designccs: Hardware Power and energyccs: Computing methodologies Machine learning
@@ 1. Introduction
Since the advent of the Exascale era, we have witnessed the convergence between High-Performance Computing (HPC) and Artificial Intelligence (AI). The ever-increasing computing power of HPC systems and their ability to manage large amounts of data made the development of more and more sophisticated machine learning (ML) techniques possible. Deep Learning (DL) is a subset of ML and uses artificial Deep Neural Networks (DNNs) with multiple layers of artificial neurons to attempt to mimic human brain behavior by learning from large amounts of data. Thanks to technological and architectural improvements, not only an increasing number of parallel high-end processors but also co-processors such as graphics processing units (GPUs) and vector/tensor computing units have been integrated into the nodes of HPC systems. This supercomputing power enabled the speed up of the automatic training phase of DNN models and their subsequent inference phase in the target application scenarios. The introduction of the pioneering AlexNet (Krizhevsky2012, ) at the ImageNet challenge in 2012 was enabled by GPU computing, outlining the value of acceleration during the training and inference phase. Since then, a multitude of DNN models have been developed for various tasks including image recognition and classification, Natural Language Processing (NLP), and Generative AI. These applications require specialized hardware accelerators, to efficiently handle the heavy computational demands of DNN algorithms. DL accelerators are currently in use in several types of computing systems spanning from ultra-low-power and resource-constrained devices on the edge up to servers, HPC infrastructures, and data centers.
@@
Scope of the survey. This survey is an attempt to provide an extensive overview of the most influential architectures to accelerate DL for high-performance applications. The survey highlights various approaches that support DL acceleration including GPU-based accelerators, Tensor Processor Units, FPGA-based accelerators, and ASIC-based accelerators, such as Neural Processing Units and co-processors on the open-hardware RISC-V architecture. The survey also includes accelerators based on emerging technologies and computing paradigms, such as 3D-stacked PIM, emerging non-volatile memories such as the Resistive switching Random Access Memory (RRAM) and the Phase Change Memory (PCM), Neuromorphic Processing Units, and Multi-Chip Modules.
@@
Overall, we have reviewed the research on DL accelerators from the past two decades, covering a significant time span of literature in this field. We have described and referenced about 230 works proposed for DL acceleration. Being DL acceleration such a prolific and rapidly evolving field, we do not claim to cover exhaustively all the research works appeared so far, but we focused on the most influential contributions. Moreover, this survey can be leveraged as a connecting point for some previous surveys on accelerators on the AI and DL field  (chen2020engineering, ; Hassanpour2022, ; gao2023acm, ; rathi2023acm, ) and other surveys focused on some more specific aspects of DL, such as the architecture-oriented optimization of sparse matrices  (reuther_hpec22, ) and the Neural Architecture Search  (Chitty2022ACMSUR, ). Another research trend in state-of-the-art AI architecture design has been moving toward transformer models. A recent survey on the full stack of optimization options on Transformer inference has been recently published in (kim2023stackoptimizationtransformerinference, ).
@@
Organization of the survey. The survey is structured in different categories and sub-categories belonging to the areas of computer architecture and hardware design. As shown in Figure 1, the proposed classification is based on several representative features of the accelerators, in order to highlight their similarities and differences. To this aim, we organized the material in a way that all research papers corresponding to multiple types of classifications are cited under each classification. For example, let us consider the work $W$, which primarily belongs to the sub-category $X$ where it makes its primary contribution. According to our classification policy, this work could be cited again in another sub-category $Y$, where it makes its secondary contribution. Moreover, under each classification, we have selectively chosen the most notable and influential works and, for each work, we focused on its innovative contributions.
The main paper is structured as follows: Section 2 provides a brief background on dominant DL topologies, while Section 3 reviews the most significant acceleration solutions based on GPUs and TPUs. Section 4 introduces four types of hardware-based accelerators: FPGA-based, ASIC-based, processing elements suitable for sparse matrices, and customized accelerators based on the open-hardware RISC-V Instruction Set Architecture (ISA). Section 5 describes DL accelerators based on emerging technologies including processing-in-memory, in-memory computing, neuromorphic accelerators, multi-chip modules, and some trends in quantum and photonic computing for deep learning. Finally, some concluding remarks on DL accelerators can be found in Section $6$. The survey is complemented by a supplementary document (referenced as Appendix A) including some basic definitions and in-depth technological aspects. To conclude, we hope this survey could be useful for a wide range of readers, including computer architects, hardware developers, HPC engineers, researchers, and technical professionals. A major effort was spent to use a clear and concise technical writing style: we hope this effort could be useful in particular to the young generations of master’s and Ph.D. students. To facilitate the reading, a list of acronyms is reported in Table 1.
@@
dir tree [Survey Organization [§ 1 Introduction ] [§ 2 DL Background ] [§ 3 GPU- & TPU-based Accelerators [GPU-based accelerators ] [TPU-based accelerators ] ] [§ 4 Hardware Accelerators [Reconfigurable Hardware Accelerators ] [ASIC-based Accelerators ] [Accelerators for Sparse Matrices ] [Accelerators based on open-hardware RISC-V ] ] ]
@@
dir tree [ [§ 5 Acceleration-based on Emerging Technologies [Processing-in-Memory ] [In-Memory Computing ][Neuromorphic Accelerators ] [Multi-Chip Modules ] [Quantum and Photonic Computing ] ] [§ 6 Conclusions ] ]
@@
MCM: Multi-Chip Module ML: Machine Learning NDP: Near Data Processing
NN: Neural Network NoC: Network on Chip PCM: Phase Change Memory
PCU: Programmable Computing Unit PIM: Processing In-Memory PULP: Parallel Ultra Low Power
QC: Quantum Computing QML: Quantum Machine Learning QNN: Quantized Neural Network
@@
RISC: Reduced Instruction Set Computer RNN: Recurrent Neural Network SoC: System on Chip
SP: Single Precision SIMD: Single Instruction Multiple Data SIMT: Single Instruction Multiple Thread
SNN: Spiking Neural Network SRAM: Static Random Access Memory TPU: Tensor Processing Unit
TNN: Ternary Neural Network VPU: Vector Processing Unit VRAM: Video Random Access Memory
@@ 2. Deep Learning Background
Deep Learning (lecun2015, ; Schmidhuber_2015, ) is a subset of ML methods that can automatically discover the representations needed for feature detection or classification from large data sets, by employing multiple layers of processing to extract progressively higher-level features. The most recent works in literature clearly show that two main DL topologies have emerged as dominant: Deep Neural Networks and Transformers.
@@ 3.1. GPU-based accelerators
Graphics Processing Units (GPUs) are specific-purpose processors introduced to compute efficiently graphics-related tasks, such as 3D rendering. They became widely used since the nineties as co-processors, working alongside Central Processing Units (CPUs) to offload graphics-related computations. The introduction of programmable shaders into GPU architectures, increased their flexibility paving the way for their adoption to perform general-purpose computations. Despite being specifically designed for computer graphics, their highly parallel architecture is well suited to tackle a wide range of applications. Consequently, in the early 2000s, GPUs started to be used to accelerate data-parallel computations not necessarily related to graphics, which could benefit from their architecture as well. This practice is commonly referred to as General-Purpose computing on Graphics Processing Units GPUs (GP-GPU) and started to be increasingly popular in the early 2010s with the advent of the CUDA language. The technological development of the last ten years significantly increased the computing power of GPU devices, which, due to their highly parallel nature, are incidentally very well suited to accelerate neural network training algorithms. The availability of such computing power allowed more complex neural network models to become practically usable, fostering the development of DNNs.
The impressive results obtainable with DNNs in the context of AI, followed by significant investments in this market sector, induced hardware manufacturers to modify GPU architectures in order to be even more optimized to compute such workloads, as an example implementing the support for lower-precision computations. This led to a de-facto co-design of GPU architectures and neural network algorithms implementations, which is nowadays significantly boosting the performance, accuracy, and energy efficiency of AI applications. A brief review on the basic features of GPU architectures able to boost the performance of HPC and Deep Learning applications is provided in Section A.2 of the Appendix A.
@@
Table 3. Selected features of the most recent GP-GPU systems
@@
Number of physical cores 132 220 32
Number of logical-cores 16896 14080 4096
@@
The comparison evidences that the higher the memory bandwidth provided to sustain DL workloads like model training, the higher the power consumption of the accelerator. Further, a huge number of parallel resources intended as physical cores are mandatory to achieve high computing performance to reduce model training time, however at the expense of reduced energy efficiency.
@@ 3.2. TPU-based accelerators
Google showcased the first TPU (Jouppi_2017, ; jouppiMotivationEvaluationFirst2018, ) at ISCA in 2017, but according to the original paper the first deployment occurred in 2015 – just three years after the “AlexNet revolution”. Their last TPU v4 implementation outperforms the previous TPU v3 by 2.1x and improves performance/Watt by 2.7x (jouppi2023isca, ). The architecture of the TPU presented in  (Jouppi_2017, ; jouppiMotivationEvaluationFirst2018, ) is centered on a large (256$\times$256) systolic array operating on signed or unsigned 8-bit integers and targeting exclusively data center inference applications; this is coupled with a large amount of on-chip SRAM for activations (24 MiB) and a high-bandwidth (30 GiB/s) dedicated path to off-chip L3 DRAM for weights. The next design iterations (TPUv2, TPUv3) (jouppiDomainspecificSupercomputerTraining2020, ) forced to move from an inference-oriented design to a more general engine tuned for both inference and training, employing the 16-bit BF16 floating-point format, more cores (2 per chip) using each one or two 4$\times$ smaller arrays than TPUv1 (128$\times$128, to reduce under-usage inefficiencies). TPUv2/v3 also introduced high-bandwidth memory support, which results in more than 20$\times$ increase in the available off-chip memory bandwidth. In 2019, Habana Labs and Intel proposed Goya and Gaudi as microarchitectures for the acceleration of inference (medina2019hotchips, ). Goya (medina2019hotchips, ) relies on PCIe 4.0 to interface to a host processor and exploits a design that uses a heterogenous approach comprising of a large General Matrix Multiply (GMM) engine, TPUs, and a large shared DDR4 memory pool. Each TPU also incorporates its local memory that can be either hardware-managed or fully software-managed, allowing the compiler to optimize the residency of data and reduce movement. Each of the individual TPUs is a VLIW design that has been optimized for AI applications and especially for training. The TPU supports mixed-precision operations including 8-bit, 16-bit, and 32-bit SIMD vector operations for both integer and floating-point. Gaudi has an enhanced version of the TPUs and uses HBM global memories rather than the DDR used in Goya, increasing the support towards bfloat16 data types and by including more operations and functionalities dedicated to training operations. While Google and Intel rely on a mixture of in-house designs and GPUs, the other main data center providers typically rely on NVIDIA GPUs to serve DL workloads. Starting from the Volta architecture (choquetteVoltaPerformanceProgrammability2018, ) and continuing with Ampere (choquetteA100DatacenterGPU2021, ) and Hopper (choquetteNVIDIAHopperH1002023, ; elsterNvidiaHopperGPU2022, ), NVIDIA has embedded inside the GPU Streaming Multiprocessors the counterpart of smaller TPUs, i.e., TensorCores.
@@
Concerning academic and research-proposed architectures, IBM Research focused on introducing techniques to reduce the precision of data formats used for training (agrawal20219, ; venkataramaniRaPiDAIAccelerator2021, ), introducing Hybrid-FP8 formats in training ASICs and tensor processors. A similar effort is performed by the authors of Cambricon-Q (zhaoCambriconQHybridArchitecture2021, ), which also introduces further improvements to exploit the statistical properties of tensors to minimize bandwidth consumption and maximize efficiency. Finally, Gemmini (gonzalez16mm106GOPS2021, ; gencGemminiEnablingSystematic2021, ) and RedMulE (tortorellaRedMulECompactFP162022a, ; tortorellaRedMuleMixedPrecisionMatrixMatrix2023, ) are efforts to introduce tensor processor hardware IPs (respectively, generated from a template and hand-tuned) that can be integrated inside System-on-Chips, similarly to what NVIDIA does with TensorCores. Further details on the basic features of TPU architectures are provided in Section A.2 of the Appendix A.

### v3

@@ Abstract.
Recent trends in deep learning (DL) have made hardware accelerators essential for various high-performance computing (HPC) applications, including image classification, computer vision, and speech recognition. This survey summarizes and classifies the most recent developments in DL accelerators, focusing on their role in meeting the performance demands of HPC applications. We explore cutting-edge approaches to DL acceleration, covering not only GPU- and TPU-based platforms but also specialized hardware such as FPGA- and ASIC-based accelerators, Neural Processing Units, open hardware RISC-V-based accelerators, and co-processors. This survey also describes accelerators leveraging emerging memory technologies and computing paradigms, including 3D-stacked Processor-In-Memory, non-volatile memories like Resistive RAM and Phase Change Memories used for in-memory computing, as well as Neuromorphic Processing Units, and Multi-Chip Module-based accelerators. Furthermore, we provide insights into emerging quantum-based accelerators and photonics. Finally, this survey categorizes the most influential architectures and technologies from recent years, offering readers a comprehensive perspective on the rapidly evolving field of deep learning acceleration.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURccs: Computer systems organization Architecturesccs: Hardware Reconfigurable logic and FPGAsccs: Hardware Emerging technologiesccs: Hardware Very large scale integration designccs: Hardware Power and energyccs: Computing methodologies Machine learning
@@ 1. Introduction
With the advent of the Exascale era, we have witnessed a growing convergence between High-Performance Computing (HPC) and Artificial Intelligence (AI). The increasing computing power of HPC systems, combined with their ability to manage vast amounts of data, has driven the development of more and more sophisticated machine learning (ML) techniques. Deep Learning (DL), a subset of ML, utilizes Deep Neural Networks (DNNs) with multiple layers of artificial neurons to mimic the human brain behavior by learning from large datasets. Thanks to advancements in technology and system architecture, HPC nodes now integrate not only an increasing number of high-end parallel processors but also specialized co-processors such as Graphics Processing Units (GPUs) and vector/tensor computing units. This supercomputing power has significantly accelerated both the training and inference phases of DNN models used in several application scenarios. The introduction of the pioneering AlexNet (Krizhevsky2012, ) model at the ImageNet challenge in 2012 marked a turning point, demonstrating the power of GPU acceleration in deep learning. Since then, numerous DNN models have been developed for various tasks including image recognition and classification, Natural Language Processing (NLP), and Generative AI. These applications demand specialized hardware accelerators, to efficiently handle the heavy computational workload of DNN algorithms. Today, DL accelerators are deployed across a wide range of computing systems spanning from ultra-low-power resource-constrained devices to high-performance servers, HPC infrastructures, and large-scale data centers.
@@
Scope of the survey. This survey is an attempt to provide an extensive overview of the most influential architectures to accelerate Deep Learning for high-performance applications. The survey highlights various approaches that support DL acceleration including GPU-based accelerators, Tensor Processor Units, FPGA-based accelerators, and ASIC-based accelerators, such as Neural Processing Units and specialized co-processors based on the open-hardware RISC-V architecture. The survey also includes accelerators based on emerging technologies and computing paradigms, such as 3D-stacked PIM, emerging non-volatile memories such as the Resistive Random Access Memory (RRAM) and the Phase Change Memory (PCM), Neuromorphic Processing Units, and Multi-Chip Modules.
@@
Overall, we have reviewed the research on DL accelerators from the past two decades, covering a significant period of literature in this field. Being DL acceleration a prolific and rapidly evolving field, we do not claim to cover exhaustively all the research works appeared so far, but we focused on the most influential contributions. Moreover, this survey can be leveraged as a connecting point for some previous surveys on AI and DL accelerators  (chen2020engineering, ; Hassanpour2022, ; gao2023acm, ; rathi2023acm, ) and other surveys focused on some more specific aspects of DL, including the architecture-oriented optimization of sparse matrices  (reuther_hpec22, ) and the Neural Architecture Search  (Chitty2022ACMSUR, ). Another research trend in state-of-the-art AI architecture design addresses transformer models. A recent survey on the full stack of optimizations on transformer inference has recently been published in (kim2023stackoptimizationtransformerinference, ).
@@
Organization of the survey. The survey is structured in different sections and sub-sections belonging to the areas of computer architecture and hardware design, as shown in Figure 1, To this aim, we organized the material in a way that all research papers corresponding to multiple types of sections are cited under each section. Moreover, for each section, we have selectively chosen the most notable and influential works and, for each work, we focused on its most innovative contributions.
To conclude, we hope this survey could be useful for a wide range of readers, including computer architects, hardware developers, HPC engineers, researchers, and technical professionals. A major effort was spent to use a clear and concise technical writing style: we hope this effort could be useful in particular to the young generations of master’s and Ph.D. students. To facilitate the reading, a list of acronyms is reported in Table 1.
@@
dir tree [Survey Organization [§ 1 Introduction ] [§ 2 DL Background ] [§ 3 GPU- & TPU-based Accelerators [GPU-based accelerators ] [TPU-based accelerators ] ] [§ 4 Hardware Accelerators [Reconfigurable Hardware Accelerators ] [ASIC-based Accelerators] ] ]
@@
dir tree [ [[Accelerators for Sparse Matrices ] [Accelerators based on open-hardware RISC-V ] ] [§ 5 Accelerators based on Emerging Technologies [Processing-in-Memory ] [In-Memory Computing ][Neuromorphic Accelerators ] [Multi-Chip Modules ] [Quantum and Photonic Computing ] ] [§ 6 Conclusions ] ]
@@
MCM: Multi-Chip Module ML: Machine Learning PCM: Phase Change Memory
PIM: Processing In-Memory QC: Quantum Computing QNN: Quantized Neural Network
@@
RISC: Reduced Instruction Set Computer RNN: Recurrent Neural Network SNN: Spiking Neural Network
SoC: System on Chip SRAM: Static Random Access Memory TPU: Tensor Processing Unit
@@ 2. Deep Learning Background
Deep Learning (lecun2015, ; Schmidhuber_2015, ) is a subset of ML methods that can automatically discover the representations needed for feature detection or classification from large data sets, by employing multiple layers of processing to extract progressively higher-level features. The most recent works in literature clearly show that two main DL topologies have emerged as dominant: DNNs and Transformers.
@@ 3.1. GPU-based accelerators
Graphics Processing Units (GPUs) are specific-purpose processors introduced to compute efficiently graphics-related tasks, such as 3D rendering. They became widely used since the nineties as co-processors, working alongside Central Processing Units (CPUs) to offload graphics-related computations. The introduction of programmable shaders into GPU architectures increased their flexibility paving the way for their adoption to perform general-purpose computations. Despite being specifically designed for computer graphics, their highly parallel architecture is well suited to tackle a wide range of applications. Consequently, in the early 2000s, GPUs started to be used to accelerate data-parallel computations not necessarily related to graphics. This practice is commonly referred to as General-Purpose computing on Graphics Processing Units GPUs (GP-GPU) and started to be increasingly popular in the early 2010s with the advent of the CUDA language. The technological development of the last ten years significantly increased the computing power of GPUs, which, due to their highly parallel nature, are incidentally very well suited to accelerate neural network training algorithms. The availability of such computing power allowed more complex neural network models to become practically usable, fostering the development of DNNs.
The impressive results obtainable with DNNs, followed by significant investments in this market sector, induced hardware manufacturers to modify GPU architectures in order to be even more optimized to compute such workloads, as an example implementing the support for lower-precision computations. This led to a de-facto co-design of GPU architectures and neural network algorithms implementations, which is nowadays significantly boosting the performance, accuracy, and energy efficiency of AI applications. The basic features of GPU architectures able to boost the performance of HPC and DL applications are briefly reviewed in Section A.2 of the Appendix A.
@@
Table 3. Selected features of the most recent GP-GPU systems.
@@
The comparison evidences that the higher the memory bandwidth provided to sustain DL workloads like model training, the higher the power consumption. Further, a huge number of parallel resources intended as physical cores are mandatory to achieve high computing performance to reduce model training time, however at the expense of reduced energy efficiency.
@@ 3.2. TPU-based accelerators
Google showcased the first TPU (Jouppi_2017, ; jouppiMotivationEvaluationFirst2018, ) at ISCA in 2017, but according to the original paper the first deployment occurred in 2015 – just three years after the “AlexNet revolution”. Their last TPU v4 implementation outperforms the previous TPU v3 by 2.1x and improves performance/Watt by 2.7x (jouppi2023isca, ). The architecture of the TPU presented in  (Jouppi_2017, ; jouppiMotivationEvaluationFirst2018, ) is centered on a large (256$\times$256) systolic array operating on 8-bit integers and targeting exclusively data center inference applications; this is coupled with a large amount of on-chip SRAM for activations (24 MiB) and a high-bandwidth (30 GiB/s) dedicated path to off-chip L3 DRAM for weights. The next design iterations (TPUv2, TPUv3) (jouppiDomainspecificSupercomputerTraining2020, ) forced to move from an inference-oriented design to a more general engine tuned for both inference and training, employing the 16-bit BF16 floating-point format, more cores (2 per chip) using each one or two 4$\times$ smaller arrays than TPUv1 (128$\times$128, to reduce under-usage inefficiencies). TPUv2/v3 also introduced high-bandwidth memory support, which results in more than 20$\times$ increase in the available off-chip memory bandwidth. In 2019, Habana Labs and Intel proposed Goya and Gaudi as microarchitectures for the acceleration of inference (medina2019hotchips, ). Goya (medina2019hotchips, ) relies on PCIe 4.0 to interface to a host processor and exploits a heterogenous design approach comprising of a large General Matrix Multiply (GEMM) engine, TPUs, and a large shared DDR4 memory pool. Each TPU also incorporates its local memory that can be either hardware-managed or fully software-managed, allowing the compiler to optimize the residency of data and reduce movement. Each of the individual TPUs is a VLIW design optimized for AI applications and especially for training. The TPU supports mixed-precision operations including 8-bit, 16-bit, and 32-bit SIMD vector operations for both integer and floating-point. Gaudi has an enhanced version of the TPUs and uses HBM global memories rather than the DDR used in Goya, increasing the support towards bfloat16 data types and including more operations and functionalities dedicated to training operations. While Google and Intel rely on a mixture of in-house designs and GPUs, the other main data center providers typically rely on NVIDIA GPUs to serve DL workloads. Starting from the Volta architecture (choquetteVoltaPerformanceProgrammability2018, ) and continuing with Ampere (choquetteA100DatacenterGPU2021, ) and Hopper (choquetteNVIDIAHopperH1002023, ; elsterNvidiaHopperGPU2022, ), NVIDIA has embedded inside the GPU Streaming Multiprocessors the counterpart of smaller TPUs, i.e., TensorCores.
@@
IBM Research focused on reducing the data precision used for training (agrawal20219, ; venkataramaniRaPiDAIAccelerator2021, ), by introducing Hybrid-FP8 formats in training ASICs and tensor processors. Further improvements were achieved with Cambricon-Q (zhaoCambriconQHybridArchitecture2021, ), which exploits the statistical properties of tensors to minimize bandwidth and maximize efficiency. Finally, Gemmini (gonzalez16mm106GOPS2021, ; gencGemminiEnablingSystematic2021, ) and RedMulE (tortorellaRedMulECompactFP162022a, ; tortorellaRedMuleMixedPrecisionMatrixMatrix2023, ) introduce tensor processor hardware IPs (respectively, generated from a template and hand-tuned) that can be integrated inside System-on-Chips, similarly to what NVIDIA does with TensorCores. Further details on TPU architectures are provided in Section A.2 of the Appendix A.